OpenStudy (anonymous):

9K^3-9m^2K ------------ is equivalent to 3k+ (?) 3K^2-3mK

OpenStudy (anonymous):

(A) m (B) 3m (C) 1 (D) 2 (E) 3

OpenStudy (mertsj):

\[9k^3-9m^2k=9k(k^2-m^2)=9k(k-m)(k+m)\]

OpenStudy (anonymous):

Do I distribute? @Mertsj

OpenStudy (mertsj):

Factor the denominator and cancel any common factors.

OpenStudy (anonymous):

3k(k^3-m) ...?

OpenStudy (mertsj):

\[3k^2-3mk=3k(k-m)\]

OpenStudy (anonymous):

how did you get rid of the ^3?

OpenStudy (mertsj):

The denominator does not have an exponent 3

OpenStudy (anonymous):

oh..

OpenStudy (mertsj):

\[\frac{9k^3-9m^2k}{3k^2-mk}=\frac{9k(k-m)(k+m)}{3k(k-m)}=3(k+m)\]

OpenStudy (anonymous):

Is it B?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thank you!